OpenStudy (anonymous):

what is the percent for 1/10.

OpenStudy (anonymous):

10%

OpenStudy (hoblos):

10%

OpenStudy (anonymous):

thanks!!!

OpenStudy (darthsid):

Think of "percent" as over 100 (per cent) So, 1 over 10 is the same as 10 over 100, hence 10 %

OpenStudy (anonymous):

Percent (%) means there is only 100 in denominator.

OpenStudy (anonymous):

\[ 10\% = 10/100= 1/10 \]